Tickets to DC'S Mr.DCHS are $3per student and $5 per adult.The number of adult tickets sold was 195 more than the number of student tickets. If DCHS collected $3943 for the tickets, how many tickets were sold to students? How many were sold to adults?
Answer by macston(5194)

.
S=student tickets; A=adult tickets=S+195
.
$3S+$5A=$3943
$3S+$5(S+195)=$3943
$3S+$5S+$975=$3943
$8S=$2968
S=371
ANSWER 1: 371 student tickets were sold.
.
A=S+195
A=371+195
A=566
ANSWER 2: 566 adult tickets were sold.
.
CHECK:
$3S+$5A=$3943
$3(371)+$5(566)=$3943
$1113+$2830=$3943
$3943=$3943
.